OpenStudy (katie327):

bethany needs to borrow $8000 . she can borrow the money at a 5.5% simple interest for 5 year or borrow at 3% with interest compounded continuously for 5 years. A. how much total interest would she pay at 5.5% simple interest? B. how much total interest at 3% compounded continuously ? c. which option results in total less interest?

OpenStudy (katie327):

I am not sure If I am doing this right. For A. I got 440 total interest For B. I got 1200 total interest.

satellite73 (satellite73):

i have no idea how to do the first one for the second, compute \[\huge 8,000\times e^{.15}\]

jimthompson5910 (jim_thompson5910):

for part a, use the formula i = P*r*t i = simple interest earned P = deposit amount r = interest rate in decimal form t = time in years

satellite73 (satellite73):

that will give you the total amount after three years subtract 8,000 to get the amount that is interest

OpenStudy (katie327):

I got A it was 2200

satellite73 (satellite73):

btw no one compounds interest continuously, that is something only found in nature

jimthompson5910 (jim_thompson5910):

`I got A it was 2200` correct